What is the index of refraction of a medium in which light travels at 1.89x10^8m/s


1
Expert's answer
2021-05-24T16:35:27-0400

n=index of refraction

c=speed of light in a vacuum "(3\times10^8\frac{m}{s})"

v=speed of light in the other medium"(\frac{m}{s})"

"n=\frac{c}{v}"

"n=\frac{3\times10^8\frac{m}{s}}{1.89\times 10^8\frac{m}{s}}"

"n=1.59"